# Epistemological Bias As a Means for the Automated Detection of Injustices in Text

### v1

@@ 3.5 Dataset source used
Lexicons used in the lexicon lookup were from social science research and collated in Pryzant et al. (2020a). We compiled these lexicons into one large dictionary and included some metadata about the lexicon i.e. source, creators, resources about the epistemological bias type, etc. Solely using an epistemological lexicon look up directly to find the biased words will be limited in its performance, for various reasons e.g. distributional shift, lexicons cannot scale well enough without requiring regular manual auditing, updating of the lexicon databases - when newer forms of subtle bias words arise, etc. There are various studies which show the limitations of using just a lexicon based approach alone, for example in Cryan et al. (2020), the authors discuss on page 8 and show in Table 5 how lexicon approaches perform less accurately than an end to end approach using BERT to detect gender based stereotypes. Therefore, it is beneficial to not solely lean on lexicons and we include it as a contributor in detecting injustices in our framework.
@@
When it comes to Meghan Markle and Kate Middleton, two people associated to the British royal family by marriage, the media depicts several aspects of their lives which they share differently and often can be reflective of framing or character injustices [^4] https://www.buzzfeednews.com/article/ellievhall/meghan-markle-kate-middleton-double-standards-royal. We use actual headlines (examples as seen in Table 3 in the Appendix) from such depictions in our comparative test to further illustrate the detection and harms of these injustices. We scale the inputs to our framework from an initial 20 article headlines used in the comparative test by gathering 1645 articles from the Bing API (up to 100 articles for each topic from no more than 1 month back). We searched the topics listed in Appendix A. The results of these comparative tests are further discussed in our results (Section 4).
4 Results and Discussions
4.1 Tagger Model Ablation Study
@@
Table 1: Ablation study results showing values for the first 3 epochs. Note: A high score is better for accuracy and a lower score is better for loss. Bold values indicate when the evaluation loss was less than the previous step.
As shown in Table 1 and Figure 3, we observe a lower training loss value between epoch 3 & 6 during the initial training using the dataset which included the TMI feature - as compared to when we do not. TMI also gives a better evaluation accuracy and evaluation loss as compared to the others before any overfitting is observed (which causes the evaluation loss to exponentially increase from about epoch 3). This overfitting occurs as a result of using too many epochs during the initial training of the tagger model. When fine-tuning a BERT model, a good practice as suggested by the authors of BERT is to use between 2 and 4 epochs. In light of this, and from the observations made from the initial training experiments, we save and use as the best point the tagger model after epoch 2.
Figure 3: Training and Validation Results for the Tagger Model using 10 epochs. no features = the expert features as mentioned in section 3.1 was not added, with expert features = combine the expert features as mentioned in section 3.1 and with expert + tmi features = combine the expert features whilst using the data samples containing TMI class no tmi or tmi
4.2 Comparative Test
We performed a comparative test to illustrate how our framework shows injustices e.g. character and framing injustice. Particularly, we capture the type of epistemological bias attached to the tagged word in each entry, observe if a relevant stereotype is associated with the inputs, and observe the depiction of different subjects. In this study, we look at Meghan Markle and Kate Middleton as our subjects of interest. Meghan is associated to the British royal family, by marriage, and resigned from her royal duties due to abuse from the media and royal family (e.g. the Firm). Kate is also a member of the British royal family, by marriage, and was promoted to Princess of Wales. Even before Meghan Markle had resigned from her royal duties, the media sought to minimize her experience and diminish her comments and character. We often see instances of Meghan as the subject of an article and the media using sarcasm and criticism towards her, thus character injustice. Such acts have led readers to having tainted images of Meghan and even not believing her statements or actions, thus testimonial injustices. Yet, for the same and similar topics of concern, media members speak charmingly about Kate, thus leading to framing injustices shown. It is important to note here, that Kate Middleton and Meghan Markle shared similar positions, interests, and abilities. With this, we see them as good subjects to observe in our comparative test.
Figure 4: Epistemological Bias Types - Comparative Test
@@ 4.2 Comparative Test
6 Meghan family law 0.3244 arab folks are not protected 0.2951
10 Kate women should wear dresses 0.3131 women should wear dresses 0.3131
20 Meghan women should dress modestly 0.3840 women should wear dresses that cover themselves 0.4047
@@
Table 2: Table showing the top ranked potential stereotype and stereotype concept (semantic distance $>0.3$ to the headline. The No. column represents the headline number from Table 3. See full list of result in Table 5 in Appendix 3.
@@
In Figure 4, we plotted results of an experiment where we initially analysed 10 article headlines of Meghan Markle and 10 articles of Kate Middleton through our framework. Each article was chosen because it discusses a topic which is common between each of the two subjects and allows us to see the differences in how the two subjects are spoken about concerning that topic. It has been acknowledged by several outlets that these particular articles which we have mentioned are unjust4. Majority of these articles are seen in the aforementioned articles to show these comparisons, but also a few additional articles were chosen by our team. In this plot, we capture the subject of the headline (Meghan/Kate), sentiment of the entire headline sentence, and the associated epistemological bias types for each headline as shown in table 4 located in the Appendix. See the complete list of headlines in Table 3 in Appendix A. The light blue circles represent the overall sentiment of the headlines, we will refer to them as the sentiment headline level. We use the 3 categories of positive, neutral, and negative sentiments. Within each sentiment headline level, we capture the subject of the headline, we will call this the subject level. Note that in Figure 4, Kate takes up the entire positive sentiment space, there were no examples that were detected to have positive sentiments for Meghan. Within each subject level, we capture the epistemological bias types associated with the tagged words in the headlines for each subject; we will refer to this as the bias-type level. This illustration shows us how much space is filled by each subject.
@@
From this plot, we can see the headline sentiment level, the negative sentiment circle is mostly filled with articles about Meghan and the most common bias type is positively subjective. The intuition here is that many articles about Meghan contain sarcasm, though this is not the only indicator. There was one article about Kate which had negative sentiment and each of the bias types in her circle were identified in the headline, this is why they are all equal in size. In looking at the headline sentiment level, the positive sentiment circle is completely filled with articles about Kate and the most common bias type is implicatives. The intuition here is that, in articles which have common topics between Meghan and Kate, only Kate is talked about in a suggestively positive way. In looking at the headline sentiment level, the neutral sentiment circle is mostly filled with articles about Kate and the most common bias type is positively subjective and regular. Recall, regular is the epistemological bias type assigned to detected words which do not appear, nor do similar words appear in the lexicon lookup. We can also observe, the articles which appear for Meghan with neutral sentiment are negatively subjective in nature, informing us they likely are not sarcastic but have negative suggestions. For Kate, she is talked about in a suggestively positive way, even in the neutral sentiments. This analysis confirms that when there is a topic that Meghan and Kate share - there was a use of words for Meghan which are sarcastic and cause framing and character injustices to her as a subject. Both of which can contribute to someone taking Meghan less seriously, thus causing testimonial injustice.
@@
4.2.1 Scaling the comparative test
@@
In Figures 8(a) and 8(b) in Appendix A, we see the plots obtained from scaling this analysis. Supporting our previous discussion, the observation from the Negative Sentiment circle is that Meghan often has more implicit subjective biases in a negative sentimental context compared to Kate. Likewise, for Kate we observed that the implicit subjective biases detected in the text associated with her are mainly used in a positively sentimental context (she takes up more space in the overall Positive Sentiment circle compared to Meghan). Most of the sample headline sentences had a Neutral Sentiment, hence the reason why the outer circle in Figure 8(b) showing the neutral sentimental context is extremely large. We observed that both subjects had almost equal amount of the epistemological bias types in this neutral context.
5 Conclusions
@@
We describe a novel framework which uses a unique combination of different NLP techniques to detect character, testimonial, and framing injustices in text. These forms of injustices are often subtle and hard to quickly detect. The framework includes a fine-tuned BERT based epistemological tagging model, a GPT based stereotype generative model and SBF model, semantic similarity searching and lexicon lookup of epistemological biased words from the social science field. We provide qualitative empirical evidence as a justification for using this framework in media settings to detect such injustices. Further, we show our proposed UI (aimed at helping editors and authors of text content), by automatically detecting and explaining potential bias types and injustices that might be present in their content. We anticipate this UI will encourage them to take necessary, preventative steps in avoiding unjust acts pre-publication.
@@
5.1 Potential issues elevated due to the proposal and means to resolve them
A critical area for bias in systems and models’ design often stem from a given human’s intrinsic biases. They are usually a reflection of ourselves. One possible solution to this problem is to ensure that a diverse group of individuals are involved from the inception of the solutions design to the testing phase of such technical solutions—this is one reason why we gathered a diverse group of authors to be involved in the discussions presented in this paper.
A second area for bias surrounds the definition of terms and assumptions biases. A way we attempt to resolve this is by considering the existing consensus definitions of epistemological bias from the social sciences, where media bias has been studied for decades.

### v2

@@ 3.5. Datasets
Lexicons used in the lexicon lookup were from social science research and collated in Pryzant et al. (2020a). We compiled these lexicons into one large dictionary and included some metadata about the lexicon i.e., source, creators, resources about the epistemic bias type, etc. Solely using an epistemic lexicon look up directly to find the biased words will be limited in its performance, for various reasons e.g., distributional shift, lexicons cannot scale well enough without requiring regular manual auditing, updating of the lexicon databases - when newer forms of subtle bias words arise, etc. There are various studies which show the limitations of using just a lexicon based approach alone, for example in Cryan et al. (2020), the authors discuss how lexicon approaches perform less accurately than an end to end approach using BERT to detect gender based stereotypes. Therefore, it is beneficial to not solely lean on lexicons and we include it as a contributor in detecting injustices in our framework.
@@
When it comes to Meghan Markle and Kate Middleton, two people associated to the British royal family by marriage, the media depicts several aspects of their lives which they share differently and often can be reflective of framing or character injustices [^4] https://www.buzzfeednews.com/article/ellievhall/meghan-markle-kate-middleton-double-standards-royal. We use actual headlines (some are shown in Table 1 with full list of examples in Table 5 of the Appendix A) from such depictions in our comparative test to further illustrate the detection and harms of these injustices. We scale the inputs to our framework from an initial 20 article headlines used in the comparative test by gathering 1645 articles from the Bing API (up to 100 articles for each topic from no more than 1 month back). We searched the following topics: “Kate Middleton”, “Meghan Markle”, “Kate Middleton family”, “Meghan Markle family”, “Kate Middleton outfit”, “Meghan Markle outfit”, “Kate Middleton child”, “Meghan Markle child”, “Kate Middleton charity”, “Meghan Markle charity”, “Kate Middleton wedding”, “Meghan Markle wedding”, “Kate Middleton moving”, “Meghan Markle moving”, “Kate Middleton bullied”, “Meghan Markle bullied”, “Kate Middleton dress”, “Meghan Markle dress”, “Kate Middleton book”, “Meghan Markle book”, “Kate Middleton makeup”, “Meghan Markle makeup”,“Kate Middleton health”, “Meghan Markle health”, “Kate Middleton photo”, “Meghan Markle photo”, “Kate Middleton money”, “Meghan Markle money”. The results of these comparative tests are further discussed in our results (Section 4).
Sentence No. Headline Subject 1 Meghan Markle spent a staggering £38,000 on her clothes for a charity trip Meghan 2 Kate Middletons £100,000 Astonishing value of the dress that won a Prince’s heart (and has hung in a wardrobe for eight years) Kate 3 Meghan Markles beloved avocado linked to human rights abuse and drought, millennial shame Meghan 4 Kates morning sickness cure? Prince William gifted with an avocado for pregnant Duchess Kate
Table 1. Example headline sentences used. For a full list of the initial 20 examples, see Appendix 5
3.6. Validation Survey
To validate our framework components, we conducted a validation survey with persons who should have some expertise in writing and/or editing. The survey included several questions aimed at evaluating the effectiveness of our framework components and pipeline to assist in automatically identifying areas for improvement for journalists and editors striving for fairness and neutrality in their text. Therefore we conduct a survey to validate the credibility and viability of our approach. We sought participants who work/volunteer in a writing/editing centers, are Senior/Grad Level journalism students, education students, educators, journalists, editors, publishers, bloggers, authors/ editors/reviews of a published work, or have operated as a reviewer for a journal/conference that accepts publications. The survey was approved by our local Institutional Review Board (IRB) and conducted through an online survey form. The respondents had unlimited time to respond and could only submit a single response. We had a total of 30 respondents, of which 8 were removed due to completing multiple surveys, improperly answering the questions (i.e., giving random answers to the survey questions), or refusal to adhere to directions of the investigator. The survey presented 16 of the headlines from Table 5 (8 pertaining to Kate, 8 pertaining to Meghan) and asked these questions about the headlines:
  1. (1)
    Main epistemic biased word is:, What alternative word or phrase will you use?
  2. (2)
    If there is an injustice present, how would you label it? (Framing Injustice, Character Injustice, Testimonial Injustice, None)
  3. (3)
    Is there sarcasm present in this headline? (yes, no)
  4. (4)
    What is the sentiment of this headline? (Positive, Neutral, Negative, Don’t know)
  5. (5)
    Which stereotype presented here is the most related to this headline?
Questions 1 and 2 were freeform, short-answer, while questions 3 - 6 were multiple-choice responses. Question 6 asked for the stereotypes based on the top 4 responses from the SBF and CO-Star Models. Lastly, to validate our assumption on adding a TMI feature, we tested these prompts which are from our TMI headline dataset and the modified versions of the headlines (containing less descriptors for the first headline and more descriptors for the remaining headlines): “Brazil: evangelical superstar expelled from Congress over alleged role in husband’s murder”, “Woman unseated from Congress due to alleged role in husband’s murder”, “Last Man Standing: Suge Knight and the Murders of Biggie and Tupac review - a rueful return to Death Row”, ‘Suge Knight and the Murders of Biggie and Tupac, former opponents and rappers, review - a shameful return to Death Row”, “Grace Millane trial concludes with New Zealand male convicted as guilty of murdering the British woman backpacker”, and “Grace Millane trial: New Zealand man found guilty of murdering British backpacker” - with the following questions:
  1. (1)
    What is the sentiment of this headline? (Positive, Neutral, Negative, Don’t know)
  2. (2)
    How do you perceive the subject of the text?
  3. (3)
    What word is framing the sentence here?
  4. (4)
    How difficult is it for you to find the framing word (1-5, 5 being most difficult)? (1,2,3,4,5)
Questions 2 and 3 were freeform, short-answer; while questions 1 and 4 were multiple-choice responses.
4. Results and Discussions
4.1. Tagger Model Ablation Study
@@
Table 2. Ablation study results showing values for the first 3 epochs. Note: A high score is better for accuracy and a lower score is better for loss. Bold values indicate when the evaluation loss was less than the previous step.
As shown in Table 2 and Figure 5, we observe a lower training loss value between epoch 3-6 during the initial training using the dataset than when we included the TMI feature. TMI also gives a better evaluation accuracy and evaluation loss as compared to the others before any overfitting is observed (which causes the evaluation loss to exponentially increase from about epoch 3). This overfitting occurs as a result of using too many epochs during the initial training of the tagger model. When fine-tuning a BERT model, a good practice as suggested by the authors of BERT is to use between 2 and 4 epochs. In light of this, and from the observations made from the initial training experiments, we save and use as the best point the tagger model after epoch 2.
Figure 5. Training and Validation Results for the Tagger Model using 10 epochs. no features = the expert features as mentioned in Section 3.1 was not added, with expert features = combine the expert features as mentioned in Section 3.1 and with expert + tmi features = combine the expert features whilst using the data samples containing TMI class no tmi or tmi
4.2. Comparative Test
We performed a comparative test to illustrate how our framework shows injustices e.g., character and framing injustice. Particularly, we capture the type of epistemic bias attached to the tagged word in each entry, observe if a relevant stereotype is associated with the inputs, and observe the depiction of different subjects. In this study, we look at Meghan Markle and Kate Middleton as our subjects of interest. Meghan is associated to the British royal family, by marriage, and resigned from her royal duties due to abuse from the media and royal family (e.g., the Firm). Kate is also a member of the British royal family, by marriage, and was promoted to Princess of Wales. Even before Meghan Markle had resigned from her royal duties, the media sought to minimize her experience and diminish her comments and character. We often see instances of Meghan as the subject of an article and the media using sarcasm and criticism towards her, thus character injustice. Such acts have led readers to having tainted images of Meghan and even not believing her statements or actions, thus testimonial injustices. Yet, for the same and similar topics of concern, media members speak charmingly about Kate, thus leading to framing injustices shown. It is important to note here, that Kate Middleton and Meghan Markle shared similar positions, interests, and abilities. With this, we see them as good subjects to observe in our comparative test.
@@ 4.2. Comparative Test
4 Kate british women are marginalized for a joke 0.2083 pregnancy 0.3918
@@
Table 3. Table showing the top ranked potential stereotype and stereotype concept (semantic distance >0.3>0.3 to the headline. The No. column represents the headline number from Table 1. See full list of result in Table 7 in Appendix 5.
Figure 6. Epistemic Bias Types - Comparative Test
@@
In Figure 6, we plotted results of an experiment where we initially analysed 10 article headlines of Meghan Markle and 10 articles of Kate Middleton through our framework. Each article was chosen because it discusses a topic which is common between each of the two subjects and allows us to see the differences in how the two subjects are spoken about concerning that topic. It has been acknowledged by several outlets that these particular articles which we have mentioned are unjust4. Majority of these articles are seen in the aforementioned articles to show these comparisons, but also a few additional articles were chosen by our team. In this plot, we capture the subject of the headline (Meghan/Kate), sentiment of the entire headline sentence, and the associated epistemic bias types for each headline as shown in table 6 located in Appendix A. See the complete list of headlines in Table 5 in Appendix A. The light blue circles represent the overall sentiment of the headlines, we will refer to them as the sentiment headline level. We use the 3 categories of positive, neutral, and negative sentiments. Within each sentiment headline level, we capture the subject of the headline, we will call this the subject level. Note that in Figure 6, Kate takes up the entire positive sentiment space, there were no examples that were detected to have positive sentiments for Meghan. Within each subject level, we capture the epistemic bias types associated with the tagged words in the headlines for each subject; we will refer to this as the bias-type level. This illustration shows us how much space is filled by each subject.
@@
From this plot, we can see the headline sentiment level, the negative sentiment circle is mostly filled with articles about Meghan and the most common bias type is positively subjective. The intuition here is that many articles about Meghan contain sarcasm, though this is not the only indicator. There was one article about Kate which had negative sentiment and each of the bias types in her circle were identified in the headline, this is why they are all equal in size. In looking at the headline sentiment level, the positive sentiment circle is completely filled with articles about Kate and the most common bias type is implicatives. The intuition here is that, in articles which have common topics between Meghan and Kate, only Kate is talked about in a suggestively positive way. In looking at the headline sentiment level, the neutral sentiment circle is mostly filled with articles about Kate and the most common bias type is positively subjective and regular. Recall, regular is the epistemic bias type assigned to detected words which do not appear, nor do similar words appear in the lexicon lookup. We can also observe, the articles which appear for Meghan with neutral sentiment are negatively subjective in nature, informing us they likely are not sarcastic but have negative suggestions. For Kate, she is talked about in a suggestively positive way, even in the neutral sentiments. This analysis confirms that when there is a topic that Meghan and Kate share - there was a use of words for Meghan which are sarcastic and cause framing and character injustices to her as a subject. Both of which can contribute to someone taking Meghan less seriously, thus causing testimonial injustice.
@@
4.2.1. Scaling the comparative test
@@
(a) Positive and Negative Sentimental Context - Scaled Comparative Test. Note the epistemic bias types are abbreviated as such for spacing concerns - Assert: Assertives, Ent: Entailments, Hed: Hedges, Fact: Factives, Neg: Negative, Pos: Positive, Sub: Subjective.
(b) Neutral Sentimental Context - Scaled Comparative Test
Figure 7. Epistemic Bias Types - Scaled Data Comparative Test.
@@
In Figures 7(a) and 7(b), we see the plots obtained from scaling this analysis. Supporting our previous discussion, the observation from the Negative Sentiment circle is that Meghan often has more implicit subjective biases in a negative sentimental context compared to Kate. Likewise, for Kate we observed that the implicit subjective biases detected in the text associated with her are mainly used in a positively sentimental context (she takes up more space in the overall Positive Sentiment circle compared to Meghan). Most of the sample headline sentences had a Neutral Sentiment, hence the reason why the outer circle in Figure 7(b) showing the neutral sentimental context is extremely large. We observed that both subjects had almost equal amount of the epistemic bias types in this neutral context.
4.3. Human Survey Discussions
Our 22 valid respondents were made up of 1 person who had operated as a reviewer for journal, 5 educators, 3 bloggers, 3 editors/publishers, 1 who had operated as a reviewer for a conference publication, 6 journalists, 1 author of a published work, and 2 PhD Students in Journalism/Education/English. We find the tagger model of the framework agrees with the biased word choice of humans 52% of the time, the stereotype 41.2% of the time (though some choices differed, still related stereotypes in 52.9% of the time), and the sentiment 66.67% of the time. So, majority of the time, the components of the framework agreed with human analysis. In Table 4, there is a short list of the top survey answers to the questions and the output of the framework (see full list in Appendix A) Figure 8. When comparing the trend of the headlines from Figure 6 and Figure 7, we see from the survey results in Figure 8 similarly that respondents were able to detect similar proportions of sentiments for each subject of the text. We do see people tended to decide more definitively that headlines were either positive or negative rather than neutral, causing those proportion of headlines to be greater. We see more diversity in the types of epistemic biases present in the neutral and positive headlines for Kate and more diversity in the types of epistemic biases present in the negative headlines for Meghan. The participants also determined that for 3 of the headlines concerning Meghan sarcasm was equally suspected and not suspected, for 2 of the 5 remaining articles about Meghan sarcasm was detected, and for 3 of the headlines concerning Kate sarcasm was detected. This might further solidify that these instances of injustice are subtle and can be difficult to detect. The main feedback from participants of the survey was it took them longer than they would like to analyze, digest, and decide on each element encompassed in the framework. Several participants indicated they had to take breaks and return to the survey while one valid non-participant indicated they wanted to participate in the survey but that it took them too much time to answer the questions, resulting in them recusing themselves from the survey. This shows us there is a need for our novel automated framework which gives similar results to the human analysis, majority of the time. Another observation from the study is that when answering Question 2 (What alternative word or phrase will you use?), many of the respondents used words which could also perpetuate some type of injustice while others simply thought it best to remove the word causing injustice all together. Two respondents who identified as a journalist and someone who had operated as a reviewer for journal, noted that though some words caused injustices, they did not want to suggest an alternative word for any headline. Another respondent who identified as an educator similarly did not give alternative word choices for several of the scenarios but found injustice. This could be due to lack of confidence for finding a better word choice or being indifferent to the injustice occurring.
When reviewing the participant’s answers for the TMI portion of the survey, participants indicated it easier for them to find the word framing the sentence and causing injustices when there were more descriptive words present. This concurs with our hypothesis mentioned in Section 4.1. They also tended to use less harsh words when describing their thoughts of the subjects of the headlines with less descriptors, even when they felt it was more neutral and less negative.
Figure 8. Comparative Test for Survey Participant Defined Tagged Words and Headline Sentiments
Headline Participant’s choices (# of participants) Framework Component’s choice (confidence)
Meghan Markle spent a staggering £38,000 on her clothes
for a charity trip
Tagged word: staggering (20)
epibias_type: subjective
sentiment: negative (15)
stereotype: personal spending habits (18)
Injustice Present: Character (14)
Alternative word: significant (3)
Tagged word: staggering (0.999498)
epibias_type: subjective
sentiment: negative
stereotype: personal spending habits
Kate Middleton’s £100,000 Astonishing value of the dress that
won a Prince’s heart (and has hung in a wardrobe for eight years)
Tagged word: astonishing (20)
epibias_type: subjective
sentiment: negative (15)
stereotype: personal spending habits (18)
Injustice Present: Character (14)
Alternative word: significant (3)
Tagged word: astonishing (0.999342)
epibias_type: positively subjective
sentiment: neutral
stereotype: women should be dressed like brides
Meghan Markle’s beloved avocado linked to human rights abuse
and drought, millennial shame
Tagged word: beloved (18)
epibias_type: positively subjective
sentiment: negative (19)
stereotype: feminist are hypocrites (11)
Injustice Present: Character (13)
Alternative word: ” ” (2)
Tagged word: beloved (0.997946)
epibias_type: positively subjective
sentiment: negative
stereotype: feminism
Kate’s morning sickness cure? Prince William gifted with
an avocado for pregnant Duchess
Tagged word: cure (8), gifted (8)
epibias_type: entailments, positively subjective
sentiment: neutral (10)
stereotype: British women are marginalized for a joke (11)
Injustice Present: Character (8)
Alternative word: given (4)
Tagged word: gifted (0.877285)
epibias_type: positively subjective
sentiment: neutral
stereotype: British women are marginalized for a joke
Table 4. Short List of Comparing Survey Results for Participants and Framework Outputs
5. Conclusions
@@
We describe a novel framework which uses a unique combination of different NLP techniques to detect character, testimonial, and framing injustices in text. These forms of injustices are often subtle and hard to quickly detect. The framework includes a fine-tuned BERT based epistemic tagging model, a GPT based stereotype generative model and SBF model, semantic similarity searching and lexicon lookup of epistemic biased words from the social science field. We provide qualitative empirical evidence as a justification for using this framework to detect such injustices in text. Further, we show our proposed UI (aimed at helping editors and authors of text content), by automatically detecting and explaining potential bias types and injustices that might be present in their content. We anticipate this UI will encourage them to take necessary, preventative steps in avoiding unjust acts pre-publication.
@@
A critical area for bias in systems and models’ design often stem from a given human’s intrinsic biases. They are usually a reflection of ourselves. One possible solution to this problem is to ensure that a diverse group of individuals are involved from the inception of the solutions design to the testing phase of such technical solutions—this is one reason why we gathered a diverse set of authors to be involved in the discussions presented in this paper. Further, we were intentional in sending our survey to venues who serve marginalized communities (e.g., writing centers at Latino serving institutions, educational institutions who have a majority of minoritized student body populations, Black graduate student associations, Asian journalist associations, etc…).
A second area for bias surrounds the definition of terms and assumptions biases. A way we attempt to resolve this is by considering the existing consensus definitions of epistemic bias from the social sciences, where media bias has been studied for decades.